OpenStudy (anonymous):

Your go-cart breaks down right before the end of a race,so you have to push it over the finish line. The go cart has a mass of 85kg. B. How much force must you apply to give the go cart an accelertion of 9.0m/s C. If you push with a force of 350N what is the acceleration of the cart? D. Later you are driving your friends go cart.if his cart has an acceleration of 5.5m/s^2 and the carts engine is applying a forward force of 720N what is the mass of his go cart ?

OpenStudy (anonymous):

You can get the three answers applying Newton's second law: \[F=ma\] in B you know m and a, and need to find F. in C you know m and F, and need a in D you know F and a, and must find the new value for m

OpenStudy (anonymous):

B. F=ma (85)(0.9)=76.5 N C.a=f/m 350/85=4.117 D.m=F/a 720/5.5 =130.9 Correct ?

OpenStudy (anonymous):

yes! escept for B, you used a=0.9m/s^2 it should be 9m/s^2 also, you could check the units for C and D, they should be m/s^2 and Kg. respectively